### v2

# MetaGS: A Meta-Learned Gaussian-Phong Model for Out-of-Distribution 3D Scene Relighting

Abstract
Out-of-distribution (OOD) 3D relighting requires novel view synthesis under unseen lighting conditions that differ significantly from the observed images. Existing relighting methods, which assume consistent light source distributions between training and testing, often degrade in OOD scenarios. We introduce MetaGS to tackle this challenge from two perspectives. First, we propose a meta-learning approach to train 3D Gaussian splatting, which explicitly promotes learning generalizable Gaussian geometries and appearance attributes across diverse lighting conditions, even with biased training data. Second, we embed fundamental physical priors from the Blinn-Phong reflection model into Gaussian splatting, which enhances the decoupling of shading components and leads to more accurate 3D scene reconstruction. Results on both synthetic and real-world datasets demonstrate the effectiveness of MetaGS in challenging OOD relighting tasks, supporting efficient point-light relighting and generalizing well to unseen environment lighting maps.
1 Introduction
3D scene relighting generates novel lighting effects that interact with the observed 3D environment, with recent advances in learning-based volume rendering offering effective solutions [20, 35, 22, 39, 23, 21, 3, 29]. A typical approach captures a substantial number of multi-view images of a scene under individual lighting conditions, then trains a NeRF or 3D Gaussian splatting (3DGS) model to generalize to new point light positions [27, 12, 8, 40, 7]. Notably, such idealized data requirements are often impractical in real-world scenarios. To simulate a more realistic capturing process, we build on prior work using a One Light At a Time (OLAT) setup [33, 13, 1], where training data is collected with a moving point light and a moving camera. Recent studies have further explored low-cost OLAT capture using a smartphone flashlight as the moving light source [5].
While most OLAT methods assume a coherent lighting distribution between training and testing, real-world scenarios often involve light sources that are randomly distributed around the scene, which may result in biased training data. This leads to an out-of-distribution (OOD) issue where test-time light sources deviate from the training distribution. This OOD relighting presents a more natural and challenging task, serving as a robust test for evaluating a model’s performance under “truly” novel lighting conditions. Compared with the standard OLAT setup, it complicates the learning of true object illumination properties and geometries due to a limited training corpus with entangled, time-varying lighting and viewing directions and, more critically, spatially biased lighting. As shown in Figure 1, existing OLAT methods struggle with the OOD relighting task, often producing unrealistic lighting effects, such as chaotic specular highlights and shadows, due to overfitting to the training samples. Also, because of the lack of generalizability, when new lighting distributions are introduced, they typically require full retraining or finetuning to adapt.
Figure 1: Preliminary results: NRHints [33] and GS3 [1] struggle with out-of-distribution light positions (as specified in Section. 2), as the blurry novel view synthesis results and erroneous normal predictions indicate. GS3 [1] does not require well-defined surface normals and does not directly output the normal image.
In this paper, we present two novel techniques to enable better OOD relighting. First, we provide a pilot study on meta-learning-based 3DGS and introduce MetaGS to mitigate overfitting and improve the generalizability of the OLAT models. This approach leverages bilevel optimization techniques to effectively address the challenges associated with limited training data and complex lighting variations. Specifically, we frame OLAT as a multi-task learning problem, treating rendering under specific lighting positions as distinct tasks. The core idea is to validate the learned Gaussian illumination properties under a given lighting condition using data from other lighting conditions. The resulting second-order gradients of 3DGS parameters can explicitly encourage the model to generalize to unseen illumination rather than overfitting to specific training samples, as is common with standard 3DGS loss functions.
Furthermore, MetaGS integrates simple yet fundamental physical priors into Gaussian splatting by incorporating a learnable Blinn-Phong reflection model [2]. This approach effectively decouples different shading components—diffuse, specular, and ambient—leading to a more physically grounded understanding of the interactions between objects and lighting in the training data. This design draws inspiration from prior research [38, 27, 12] indicating that modular and compositional representations enhance generalization capabilities. The decoupled rendering improves the model’s ability to generalize to novel lighting conditions, allowing it to independently adjust each shading component according to the specific lighting scenario.
We evaluate MetaGS on both synthetic and real-world datasets. It supports efficient point-light relighting under highly constrained training illuminations and significantly outperforms existing OLAT methods in OOD relighting tasks. We further demonstrate that both meta-learning and the differentiable Phong model individually contribute to improved generalization. Notably, MetaGS generalizes well to unseen environment maps, despite being trained exclusively in OLAT scenarios.
2 Preliminaries
Definition and challenges of OLAT.
OLAT refers to a specialized setting for 3D reconstruction and relighting. This involves illuminating the subject with a single point light in sequential exposures, capturing one image at each light position. Each capture is represented as a tuple $\{O_{t},V_{t},P_{t}\}$, where $O_{t}$ denotes the observed image, $V_{t}$ represents the camera configuration, and $P_{t}$ contains lighting information. In our context, $P_{t}$ specifies the position of the point light. Unlike the “multiple lights, multiple cameras” setup, which requires multi-view images captured by multiple time-synchronized cameras for each lighting condition, OLAT relies on a single camera. This significantly simplifies data collection, particularly in scenarios with dynamic lighting. The rapidly changing light positions in OLAT introduce a new challenge for 3D reconstruction. The color of a 3D point can vary significantly, increasing ambiguity for estimating true object geometries and illumination properties.
Preliminary findings.
We summarize existing OLAT relighting methods in Table LABEL:tab:olat_methods. In the preliminary experiments, we evaluate NRHints [33], the prior art in NeRF-based relighting models. We observe that it underperforms in testing scenarios with out-of-distribution (OOD) lighting positions, a feature that we believe should be crucial for relighting methods. Specifically, in Figure 1, the lighting in the training set is arranged on one side of the hemisphere, while the lighting in the test set is arranged on the opposite side (cameras are located on both sides). In such cases, NRHints fails to generate reasonable rendering results, likely due to its implicit modeling of shadows and specular reflections. We additionally evaluate a concurrent 3DGS-based approach [1] and observe similar degraded results under the OOD relighting setup. These findings highlight that understanding intrinsic illumination properties under arbitrary lighting variations remains a challenging task, as the model tends to overfit to perspective-constrained observations and fail to leverage enough physical principles when dealing with unseen lighting distribution.
Table 1: Comparison of the OLAT relighting methods.
[34] [5] [6] [33] [1] [11] Ours
Point light
Shadow computation $\times$ $\times$
OOD light $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
3 Method
In this section, we present the details of MetaGS for addressing the OOD OLAT learning challenge:
  • Sec. 3.1: We incorporate physical shading priors within Gaussian splatting via a differentiable Phong reflection model to decouple mixed illumination components.
  • Sec. 3.2: We introduce a bilevel optimization scheme to estimate light-independent scene geometries and intrinsic illumination properties, marking an early effort in volume rendering.
  • Sec. 3.3: We discuss the entire training pipeline and the implementation details.
3.1 Differentiable Phong Model
Our MetaGS method leverages simple yet generalizable physical priors from the Blinn-Phong model [18], which captures three fundamental components of light transport: ambient, diffuse, and specular reflections. The core idea is to disentangle these illumination components by learning the interactions between (i) the normal vectors of the Gaussian points, (ii) the viewing directions, and (iii) the ray direction from the point light.
Specifically, the ambient component represents constant environmental illumination, simulating indirect scattering from surrounding surfaces to establish a baseline brightness level. The diffuse component, based on Lambertian law, describes light scattering in multiple directions from rough surfaces. The specular reflection is computed based on the angle between the light direction and the bisector ($\mathbf{h}$) between the viewing direction ($\mathbf{v}$) and the light direction ($\mathbf{l}$), with a shininess exponent that represents different degrees of glossiness.
In MetaGS, we extend the original 3DGS by introducing the decoupled computation for different shading components. As shown in Figure 2, in addition to the basic Gaussian attributes (e.g., position $\mathbf{x}$, rotation $R$, scale $S$, opacity $\alpha$, and spherical harmonic coefficients $f$), each Gaussian point is further associated with a normal vector $\mathbf{n}$, a $3$-channel diffuse color ${k}_{d}$, and a $1$-channel specular coefficient $k_{s}$. The newly added attributes enhance the model’s understanding of lighting effects, facilitating the learning process of light-independent object geometries. The overall color of a Gaussian point is determined by:
$L_{p}=L_{a}+L_{d}+L_{s}=L_{a}+\displaystyle\sum\limits_{\lvert\text{lights}\lvert}(k_{d}I_{d}+k_{s}I_{s}),$ (1)
[…]
$I_{d}=\frac{I}{r^{2}}\max(0,\mathbf{n}\cdot\mathbf{l}),\quad I_{s}=\frac{I}{r^{2}}\max(0,\mathbf{n}\cdot\mathbf{h})^{p},$ (2)
where $I$ denotes the light emitted intensity, which is a global learnable parameter, $\mathbf{l}$ denotes the point-to-light normalized vector, and $\mathbf{h}=\frac{\mathbf{v}+\mathbf{l}}{||\mathbf{v}+\mathbf{l}||}$ denotes the bisector of the point-to-camera normalized vector $\mathbf{v}$ and $\mathbf{l}$. We model the coefficients $k_{d}$ and $k_{s}$ implicitly, where $k_{s}$ is multiplied by the RGB color of the point light.
Shadow computation.
We calculate shadow effects using a BVH-based ray tracing method [7]. Similar to the camera-to-point accumulated transmittance, the received light intensity $T_{i}^{\text{light}}$ represents the total transmittance of light along the light-to-point ray. It is therefore affected by the opacity of the surfaces encountered along this path. For each point, we determine light visibility by tracing a ray from the Gaussian’s center to the light source. Since only the diffuse and specular colors are influenced by incident light intensity, we update the color of each point by incorporating the light visibility factor into the diffuse and specular terms from Eq. (1):
$L_{p}=L_{a}+\ T_{i}^{\text{light}}\sum\limits_{\lvert\text{lights}\lvert}(k_{d}I_{d}+k_{s}I_{s}).$ (3)
This explicit formulation accounts for physics-based shadowing effects, provides strong interpretability, and improves generalizability compared to the implicit, high-dimensional shadow modeling methods. However, integrating this module directly into MetaGS’s pipeline presents challenges. Under OLAT conditions, learning coherent geometry is challenging, leading to difficulties in accurately separating shadows from the scene. This, in turn, results in erroneous color predictions. The following meta-learning scheme promotes the mutual learning of scene geometry and appearance, alleviating the difficulties in learning coherent attributes under varying illuminations.
Figure 2: The model design of MetaGS: Our model decomposes the illumination effects by interacting the learned Gaussian points with rays originating from both the viewer and the light source.
3.2 Meta-Learned Gaussian Relighting
Existing 3D relighting methods exhibit performance degradation when handling out-of-distribution relighting, primarily due to overfitting lighting patterns to perspective-constrained observations, resulting in producing unreasonable lighting components (such as wrong specular and shadows).
To address this, we introduce a meta-learning framework based on bilevel optimization, which has been shown to effectively bridge the distribution shift between the training and testing domains, facilitating the generalization of optimized variables to unseen scenarios [4]. This training strategy can also promote coherent geometry learning in complex OLAT tasks, as shown in the ablation study. In MetaGS, the intuition of incorporating bilevel gradient update is to mitigate overfitting to specific light conditions by explicitly simulating test samples with OOD light sources during each gradient update, thereby improving the model’s ability to generalize to varied lighting scenarios.
As illustrated in Alg. 1, we organize the training processes with different $\{O_{t},V_{t},P_{t}\}$ as multiple learning tasks. We alternate model training between these tasks, validating the optimized variables under one lighting condition using data sampled from other conditions. This learning procedure encourages both the lighting attributes ($f,k_{a},k_{s}$) and the geometric attributes ($\mathbf{x},\mathbf{n},R,S,\alpha$) of the Gaussian points to converge cohesively.
To simulate test-time conditions, the training data is divided into support (training) and query (validation) sets. At each iteration, $2m$ samples are drawn from the training data to form the support set $\mathcal{D}_{1:m}^{\text{sup}}$ and query set $\mathcal{D}_{1:m}^{\text{query}}$ (Line 5). Each support sample pairs with a query sample, denoted by subscript $i$. During each training step, the model alternates between these tasks in the inner optimization loop (Lines 6-9) and the outer loop (Line 10).
Inner optimization loop.
In the inner loop, the model is trained on each support set to independently generate $m$ sub-models, representing the hypothetical estimates $\theta_{i}^{\prime}$ under each lighting condition. All sub-models start with the same parameters of $\theta$. The light intensity $I$ is a global parameter that is also learnable and optimized jointly with the Gaussian attributes. It is omitted here for clarity. We specify the loss function $\mathcal{L}_{\text{final}}$ in subsequent Sec. 3.3.
Outer optimization loop.
In the outer loop, a global gradient update is performed across all sampled tasks, updating the Gaussian attributes. We first compute the loss function for each query sample, $\mathcal{L}_{\text{final}}(\cdot,\cdot;\mathcal{D}_{i}^{\text{sup}})$, using the corresponding inner-loop model hypotheses $\theta_{i}^{\prime}$.We then aggregate the $m$ losses to update the initial $\theta$.This approach, involving task-specific adaptation followed by a global update, enables the model to learn generalizable representations across varying lighting conditions.
Algorithm 1 Meta-Training for OOD Relighting
[…]
2:Hyperparameters: Learning rates $\alpha,\beta$
3:Initialized parameters: Pretrained Gaussian attributes $\{\theta_{k}\}$, where $\theta_{k}=(\mathbf{x},\mathbf{n},R,S,\alpha,f_{0},k_{d},k_{s})$
4:while not converge do
5:    Sample disjoint data $\mathcal{D}_{1:m}^{\text{sup}}$ and $\mathcal{D}_{1:m}^{\text{query}}$ from $\{\mathcal{D}_{t}\}_{1:T}$
6:    for task $i$ in $\{1:m\}$ do $\triangleright$ Inner optimization loop
7:         $\theta_{i}^{\prime}\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\text{final}}(\theta;\mathcal{D}_{i}^{\text{sup}})$
8:    end for
9:    $\{\theta\}\leftarrow\{\theta\}-\beta\sum_{i=1}^{m}\nabla_{\theta}\mathcal{L}_{\text{final}}(\theta_{i}^{\prime};\mathcal{D}_{i}^{\text{query}})$triangleright Outer optimization step
10:end while
3.3 Entire Training Pipeline
We employ a three-stage training scheme. The first stage, similar to 3DGS, focuses on training the core Gaussian attributes, including position $\mathbf{x}$, rotation $R$, scale $S$, opacity $\alpha$, and spherical harmonic coefficients $f$. After this stage, the model tends to capture a rough geometry and an average color across different illuminations, serving as an effective initialization for further refinement. In the second stage, we incorporate the normal attributes into the optimization process, acknowledging that learning a Phong model heavily relies on accurate normal estimations. In the final stage, we integrate the diffuse and specular components, training all parameters concurrently through meta-learning.
Objective functions.
We follow 3DGS [9] to compute the deviation of the predicted image from the ground truth, using an L1 regularization and a D-SSIM term as the RGB loss. We also employ a sparse loss [8] to encourage the opacity values $\alpha$ of the Gaussian spheres to approach either 00 or $1$, thereby facilitating learning with opaque objects. For normal estimation, we follow GaussianShader [8] to compute the normal vectors using Gaussian’s shortest axis direction, and progressively align them with the depth-inferred pseudo-normals.
Implementation details.
We implement MetaGS using PyTorch [17]. For optimization, we use the Adam optimizer [10] with the same parameters as those specified by [9]. A consistent set of hyperparameters is applied across all scenes, with details provided in the Appendix. Our training procedure contains three stages: the first two stages each last for $4k$ iterations, followed by the final meta-learning stage of $5k$ iterations. The entire training process requires roughly one hour on a single NVIDIA RTX 3090 GPU.
4 Experiments
Figure 3: OOD relighting results on synthetic data: We present rendered novel views and error maps. While baselines often misrepresent shadows or light-dependent effects (e.g., incorrect shadows in Plastic Cup), our model better infers surface appearance.
4.1 Experimental Settings
We perform three types of OLAT relighting experiments:
  • OOD relighting: We introduce a novel view synthesis setup to assess the generalizability of 3D relighting models under OOD lighting conditions. The point lights in the training set are positioned on one side of the upper hemisphere, while the lighting in the test set is placed on the opposite side.
  • Camera-light-colocated relighting: This setup mirrors the one used in IRON [34], where the camera and point light are co-located for each image. Similar to the OOD setting, the colocated configuration also represents a constrained lighting regime, as the light source is fixed to the camera pose. For fair comparison, we generate colocated versions of the synthetic scenes accordingly.
  • Environment map relighting: We evaluate the generalization of the models exclusively trained in the OLAT setup to novel view synthesis with unseen environment lighting maps.
Figure 4: OOD relighting results on real-world data: As highlighted with the red boxes, baseline models struggle with some level of global shading consistency, such as color shifts, wrong shadows, and floating artifacts. Our approach presents physically plausible specular highlights and geometrically consistent shadows that closely match ground truths.
Table 2: PSNR results for OOD relighting. See our Appendix for full comparisons.
Method Synthetic Real-world
Ball PlasCup RubCup Cat Catsmall CupFabric Fish FurScene Pikachu Pixiu
NRHints [33] 17.25 23.92 27.44 18.04 24.63 24.65 22.57 21.55 24.00 23.03
WildLight [5] 21.73 20.95 24.02 18.65 22.53 24.03 21.47 20.33 19.09 20.22
GS3 [1] 18.84 20.30 24.37 17.66 23.34 25.04 21.12 17.34 24.11 19.63
Ours 26.76 27.54 27.95 26.45 26.44 27.29 24.68 24.82 25.54 25.65
We evaluate MetaGS on $3$ synthetic scenes and $7$ real captured scenes from NRHints [33], featuring a diverse range of materials, complex object shapes, significant self-occlusion, and intricate shadow effects. Each image is rendered (or captured) with a unique camera pose and point-light position. Full details of the datasets are provided in the Appendix.
For each synthetic scene, we generate a total of $600$ OLAT images using Blender, with $500$ for training and $100$ for testing, following out-of-distribution or colocated data patterns. For the real-world scenes, we split the data according to their point light positions to construct out-of-distribution datasets, and use 600 training images per scene—a significantly smaller subset than the full NRHints dataset.
We compare MetaGS against both NeRF-based and Gaussian-based models, with a particular focus on the OOD relighting setup. The baseline models include state-of-the-art approaches from the past two years: GS3 [1], NRHints [33], WildLight [5], Relightable 3DGS  [7], GaussianShader  [8], and IRON [34]. The evaluation metrics include PSNR, SSIM [24], and LPIPS [37].
Figure 5: Ablation studies with qualitative results: As shown in the decoupling components, the meta-learning scheme is essential for learning object illumination properties and geometries. Without it, the model generates plain lighting effects, particularly for the diffuse and specular components.
4.2 Results for Out-of-distribution OLAT Relighting
We evaluate MetaGS in scenarios with OOD point light positions to explore its generalizability to unseen illuminations. The qualitative and quantitative results are respectively presented in Figure 3, Figure 4, and Table 2. Our model remarkably outperforms the baselines in handling complex light interactions. For example, in Figure 4, the lighting and shading effects on the cat’s fur are rendered with high fidelity, while baseline models tend to produce incorrect color tones. Furthermore, as indicated by the error maps in Figure 3, all of three OLAT baselines tend to provide unrealistic results, especially for regions with significant specular reflections or shadows. Please see our Appendix for video visualizations on OOD relighting, where baselines’ performance drops as the light enters the OOD region, while our model provides better relighting results.
The generalizability of MetaGS to OOD illumination stems from two key factors. First, it is due to the Phong model’s ability to capture the underlying principles of light interaction. For example, instead of simply viewing specular highlights as the view-dependent color of a point (which is not generalizable), our method explicitly calculates the interactions of rays and surface normals, enabling the model to generate accurate highlights even in areas not directly illuminated during training. Second, meta-learning further enhances generalization by simulating test-time conditions during training, effectively reducing overfitting to specific illuminations. In the Appendix, we empirically show that the benefit of meta-learning does not come from a larger batch size.
Ablation studies.
Table 3: Ablation studies of each model component for OOD relighting. We show the average results of all three synthetic scenes.
Method PSNR SSIM LPIPS
Full model 27.42 0.9546 0.0505
w/o Meta-learning 19.14 0.8781 0.0892
w/o Shadow 21.53 0.9105 0.0735
We conduct a series of ablation studies to validate the effectiveness of the meta-learning scheme as well as the shadow computing. From Table 3 and Figure 5, we observe that the proposed meta-learning training scheme strongly impacts performance. Without it, the model struggles to converge smoothly and make unrealistic component estimations, resulting in plain rendering quality.
4.3 Generalization to Other Relighting Setups with Constrained OLAT Training
Figure 6: Relighting with environment maps: Our method, trained under OLAT settings, successfully generalizes to unseen environmental lighting; while the compared method is trained in an all-light-on setup that provides multiple lighting conditions per training view, it still fails to generalize to perform relighting with novel environment maps and exhibits visible artifacts.
Camera-light-colocated relighting.
Table 4: Novel view synthesis results in PSNR under camera-light-colocated relighting setup. Full results in all evaluation metrics are included in the Appendix.
Method Ball PlaCup RubCup
IRON [34] 26.99 34.43 36.22
Ours 38.72 36.90 38.89
We evaluate MetaGS in the camera-light-colocated setup, as introduced by IRON [34]. Like the OOD scenario, the colocated setup also imposes limited lighting diversity, as the point light is tied to the camera’s position throughout training. Quantitative results in Table 4 show that MetaGS outperforms IRON, achieving more accurate illumination inference under new point positions, and generating fewer rendering artifacts. Visualizations corresponding to these results can be found in the Appendix.
Environment map generalization.
Our model generalizes effectively to environment map relighting. We approximate global illumination by importance sampling point lights from the environment map based on their pixel intensities. Notably, for our model in particular (unlike the compared methods), the sampled light directions at test time may follow a distribution different from those encountered during training, presenting a significant challenge for OOD generalization. Instead of assuming infinite light source distance, we simulate distant lighting by placing point lights along environment directions at a fixed distance (twice the scene radius). We compare our method with R3G [7] and GaussianShader [8]. Both methods fail to converge when trained directly on out-of-distribution (OOD) data, so we construct training data that covers the full hemispherical illumination space. Even with this setup, GaussianShader still fails to converge, while Relightable 3DGS struggles with the OLAT learning task and exhibits noticeable artifacts. As shown in Figure 6, MetaGS produces diverse and visually plausible relighting results under complex environment maps, demonstrating its robustness and generalization capability.
More results.
To validate MetaGS’s capability in handling complex lighting conditions, we conduct experiments with test sets containing 22223333 light sources. We also present the results for free-viewpoint relighting with in-distribution point light positions. Please refer to the Appendix for details.
5 Related Work
Recent differentiable volume rendering techniques, including NeRF-based [16], SDF-based [31, 32], and 3DGS-based [9] methods, have significantly improved the quality and efficiency of novel view synthesis for 3D scenes. NeRF-based methods utilize deep neural networks to model volumetric scene functions, encoding color and density to synthesize high-quality images from sparse viewpoints [16, 14, 36]. However, these approaches typically require substantial training time. 3DGS significantly reduces both training and inference time. It converts point cloud data into a continuous volumetric representation by applying Gaussian kernels to point cloud data, facilitating rendering and further processing. This technique is now widely adopted for efficient 3D and 4D reconstructions across varied data types [27, 8, 15, 25, 30].
The task of 3D relighting involves altering the illumination in a 3D scene while maintaining its geometry. It requires decomposing materials and lighting of a scene from multiple images, which is challenging due to its high-dimensional nature. Recent advances in volume rendering have introduced various solutions for this task, including those based on neural fields [38, 26, 21, 19, 20, 12, 22, 21, 29] and the methods based on Gaussian point clouds [27, 8, 7, 35]. A limitation of these methods is that they require a substantial volume of multi-view images captured under individual lighting conditions, making them impractical in real-world scenarios. Several studies have proposed NeRF-based 3D relighting techniques within the OLAT framework [28, 38, 33], which greatly reduce the demands on training data. Nonetheless, these techniques tend to be computationally intensive because of NeRF’s inherent complexity in the volume rendering process. In parallel, concurrent 3DGS-based approaches [6, 1] have also explored the OLAT relighting challenge. Despite their efforts, these approaches still face difficulties when dealing with lighting conditions not encountered during training, similar to the limitations observed in NeRF-based methods.
In contrast, our method addresses this issue by incorporating specific physical priors, represented by the Phong reflection model, into the Gaussian splatting framework. We have also designed a meta-learning approach to enhance the method’s generalizability to OOD relighting scenarios.
6 Conclusions and Limitations
In this paper, we explored a novel and challenging problem: out-of-distribution (OOD) 3D relighting. To address this, we proposed MetaGS, which builds on Gaussian splatting and presented a novel bilevel optimization-based meta-learning framework that explicitly promotes generalizable Gaussian geometry and appearance learning. This meta-learning formulation allows the model to adapt to varying light sources and viewpoints, even when the training data is biased or sparsely sampled in the lighting space. Furthermore, MetaGS incorporates a differentiable Blinn-Phong reflection model within Gaussian splatting, effectively disentangling lighting effects into ambient, diffuse, and specular shading components, thereby improving the physical realism and reconstruction fidelity under diverse lighting. Extensive experiments on both synthetic and real-world datasets demonstrate that MetaGS significantly outperforms existing OLAT relighting approaches under OOD relighting conditions. It also achieves strong generalization to complex environment maps.
Despite these advances, several limitations remain. An open challenge is ensuring the robustness of our approach under more complex lighting conditions. Moreover, our current framework only accounts for direct illumination; incorporating indirect lighting could further enhance the quality of relighting. Additionally, the current model assumes simple Phong reflection, which may limit fidelity when modeling materials with strong subsurface scattering or anisotropic reflectance.
Appendix A Notations
The symbols and notations are illustrated in Table 5.
Table 5: Notations in the manuscript. Per-point variables unless being marked as global.
      Location       Notations       Meaning
      Phong model       $I$       learnable light intensity (global)
      $T_{i}^{light}$       received light intensity
      $k_{d},k_{s}$       diffuse/specular coefficient
      $I_{d},I_{s}$       diffuse/specular intensity
      $r$       point-to-light distance
      $\mathbf{n}$       normalized normal vector
      $\mathbf{v}$       point-to-camera normalized vector
      $\mathbf{l}$       point-to-light normalized vector
      Meta-learning       $\theta$       Gaussian attributes
      $\theta_{i}^{\prime}$       i-th task-specific predictor
Appendix B Datasets
We test on three synthetic scenes and seven real-world scenes from NRHints [33], as shown in Figure 7. Each image is rendered (or captured) with a unique camera pose and point-light position. For each synthetic scene, we generate a total of $600$ OLAT images using Blender, with $500$ for training and $100$ for testing, following either out-of-distribution or colocated data patterns to align with their settings. Specifically, the camera and light source are independently sampled on the upper hemisphere with centered around the scene, with elevation angles ranging from $10^{\circ}$ to $80^{\circ}$, and radial distances uniformly sampled between $4.0$× and $5.0$× the object bounding size. This setup ensures a wide diversity of incident angles while maintaining plausible visibility. In the OOD configuration, training and testing light directions are drawn from non-overlapping hemispheres (azimuth $\phi\in[0,\pi]$ for training; $\phi\in[\pi,2\pi]$ for testing) to induce generalization difficulty. In the colocated setting, camera and light positions are identical, emulating a mobile-light capture scenario. All images are rendered at a resolution of $512\times 512$ using Cycles path tracing in Blender, with $256$ samples per pixel. For real captured scenes in the out-of-distribution experiment, we manually divide the data into training and testing sets, each with lights positioned on opposite sides of the upper hemisphere.
Figure 7: Synthetic and real-world scenes used in the experiment with ground-truth images (right) and rendering results of MetaGS (left).
Table 6: Full comparison of the OOD relighting comparison on NRHints synthetic datasets.
Method Layered Woven Ball Plastic Cup Rubber Cup
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
NRHints [33] 17.25 0.8755 0.0872 23.92 0.9485 0.0470 27.44 0.9465 0.0691
WildLight [5] 21.73 0.9217 0.0568 20.95 0.9249 0.0629 24.02 0.9221 0.0789
GS3 [1] 18.84 0.9063 0.0589 20.30 0.9141 0.0628 24.37 0.9349 0.0715
Ours 26.76 0.9561 0.0423 27.54 0.9580 0.0405 27.95 0.9497 0.0688
Table 7: Full comparison of OOD relighting results on NRHints real-world datasets.
Method Cat Catsmall CupFabric Fish
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
NRHints [33] 18.04 0.7821 0.2405 24.63 0.9289 0.1094 24.65 0.9377 0.0965 22.57 0.8658 0.1396
WildLight [5] 18.65 0.7293 0.2685 22.53 0.8820 0.1448 24.03 0.9246 0.0969 21.47 0.8419 0.1458
GS3 [1] 17.66 0.6919 0.2431 23.34 0.9182 0.1097 25.04 0.9408 0.0954 21.12 0.8264 0.1401
Ours 26.45 0.8514 0.1759 26.44 0.9505 0.0980 27.29 0.9490 0.0918 24.68 0.8706 0.1386
Method FurScene Pikachu Pixiu Average
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
NRHints [33] 21.55 0.8568 0.1435 24.00 0.9279 0.0958 23.03 0.8970 0.1203 22.64 0.8853 0.1351
WildLight [5] 20.33 0.8304 0.1635 19.09 0.8941 0.1129 20.22 0.8330 0.1621 20.90 0.8479 0.1564
GS3 [1] 17.34 0.7684 0.1669 24.11 0.9203 0.0927 19.63 0.8391 0.1329 21.18 0.8436 0.1387
Ours 24.82 0.8850 0.1297 25.54 0.9350 0.0928 25.65 0.9194 0.1105 25.83 0.9087 0.1196
Appendix C More Experimental Analyses
C.1 Full Comparisons for OOD Relighting
We present the OOD relighting results with all three metrics in Table 6 and Table 7.
C.2 Clarification on Meta-learning
In our OLAT method, meta-learning helps mitigate overfitting to specific illuminations by explicitly simulating test conditions during training. Consider a renderer $R(\theta,l_{i})$, where $\theta$ represents light-independent parameters (material/geometry), and $l_{i}$ denotes the light position. Through bilevel gradients, meta-learning assesses how well the learned $\theta^{\prime}$ under $L=\|R(\theta,l_{i})-I_{i}\|^{2}$ adapts to minimizing $\|R(\theta^{\prime},l_{j})-I_{j}\|^{2}$ with the simulated test condition $l_{j}$, thereby enforcing cross-illumination consistency in geometry reconstruction. This new loss function encourages the model to learn to generalize to unseen illumination rather than to overfit to a specific training sample as direct RGB loss does. By forcing geometry parameters $\theta$ to remain consistent across $\{l_{i}^{\prime}\}$, we prevent illumination artifacts from being baked into textures, thereby achieving more accurate highlights and smoother surfaces. As shown in Table 8, increasing batch size (by task number) improves PSNR by $2+$, while meta-learning achieves an $8+$ improvement, confirming that the impact of meta-learning is not due to a larger batch size.
Table 8: A comparison of meta-learning vs. Increased batch size.
      Meta-learn       w/o Meta-Learn       Larger batch (×$\times m$)
      PSNR       27.42       19.14       20.94
C.3 Multi-light Relighting
To validate MetaGS’s capability in handling complex lighting conditions, we conduct experiments with test sets containing 22223333 light sources. As shown in Figure 8, our method demonstrates robust generalization beyond single-light scenarios.
Figure 8: Results with multiple lights (PSNR/SSIM/LPIPS).
C.4 Camera-Light-Colocated Relighting
We present the colocated relighting results with all three metrics in Table 9 and the qualitative results in Figure 9. As shown in the figure, IRON fails to accurately infer the geometry of the Layered Woven Ball scene, resulting in blurry results. In the Plastic Cup and the Rubber Cup scenes, while IRON can provide reasonable geometry reconstruction, it shows more artifacts in the appearance modeling, such as the dent of the dice. On the other hand, MetaGS can provide more accurate illumination details.
Table 9: Novel view synthesis results in the camera-light-colocated relighting setup.
Method Layered Woven Ball Plastic Cup Rubber Cup
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
IRON [34] 26.99 0.9231 0.0896 34.43 0.9808 0.0296 36.22 0.9677 0.0476
Ours 38.72 0.9904 0.0124 36.90 0.9893 0.0145 38.89 0.9712 0.0385
Figure 9: Novel view synthesis results under the camera-light-colocated relighting setup. Compared to IRON, our model more accurately infers object geometry in the Layered Woven Ball scene and produces fewer artifacts in the Plastic Cup scene.
C.5 In-distribution Novel View Relighting
This experiment follows the setup in NRHints [33], where both training and testing lights are randomly located in the upper hemisphere surrounding the scene). Each view is illuminated by a single point light placed at a unique position, though all light positions are distributed within the same region. During the testing phase, we evaluate on novel viewpoints under new, in-distribution light positions. Similar to the OOD setting, we generate $600$ OLAT images per synthetic scene using Blender ($500$ for training and $100$ for testing). For real-world scenes, we use only $600$ training images–—a significantly smaller subset compared to the full NRHints dataset.
We present the quantitative results for free-viewpoint relighting with in-distribution point light positions in Table 10. Notably, MetaGS outperforms other near real-time methods, including Relightable 3DGS and GaussianShader, by large margins, and shows an advantage over GS3, a Gaussian-based OLAT method. Our method underperforms NRHints on certain “not-so-difficult” datasets such as PlasticCup and RubberCup. This is partly because the implicit rendering model in NRHints provides better color fitting, resulting in superior in-distribution performance. While the Phong-based rendering model used in our work facilitates geometry-shading decoupling and is particularly effective for objects with incoherent geometries, such as the Woven Ball. However, does NRHints truly learn the intrinsic reflection mechanisms from the observed images? Likely not, as evidenced in the previous OOD results.
Table 10: Novel view synthesis results in the in-distribution relighting setup. We present the best and second-best results in bold and underlined format, respectively.
Method Layered Woven Ball Plastic Cup Rubber Cup Average
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
3DGS [9] 19.55 0.9130 0.0580 21.13 0.9417 0.0470 23.22 0.9248 0.0796 21.30 0.9265 0.0615
Relightable 3DGS [7] 19.62 0.9088 0.0610 21.07 0.9390 0.0502 23.20 0.9224 0.0838 21.30 0.9234 0.0650
GaussianShader [8] 19.62 0.9137 0.0586 20.92 0.9394 0.0500 23.12 0.9241 0.0833 21.22 0.9257 0.0640
NRHints [33] 18.99 0.9087 0.0731 35.81 0.9847 0.0281 36.64 0.9637 0.0481 30.48 0.9524 0.0498
WildLight [5] 22.71 0.9284 0.0499 23.28 0.9068 0.1018 24.65 0.9271 0.0737 23.55 0.9208 0.0751
GS3 [1] 28.20 0.9659 0.0396 29.53 0.9755 0.0317 29.34 0.9520 0.0618 29.02 0.9645 0.0444
Ours 30.88 0.9690 0.0281 32.51 0.9814 0.0295 31.75 0.9577 0.0570 31.71 0.9694 0.0382
Appendix D Preliminaries
D.1 3D Gaussian Splatting
Our method builds upon 3D Gaussian Splatting (3DGS) [9], which employs explicit Gaussian points for 3D scene representation. Each Gaussian’s geometry is characterized by an opacity $o\in[0,1]$, a center point $\mu\in\mathbb{R}^{3\times 1}$ and a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$:
$G(\mathbf{X})=e^{-\frac{1}{2}(\mathbf{X}-\mu)^{T}\Sigma^{-1}(\mathbf{X}-\mu)},$ (4)
where the covariance matrix $\Sigma$ can be decomposed into a scaling matrix $S\in\mathbb{R}^{3\times 1}$ and a quaternion-parameterized rotation matrix $R\in\mathbb{R}^{3\times 3}$ for differentiable optimization:
$\Sigma=RSS^{T}R^{T}.$ (5)
To render an image from a given viewpoint, 3DGS uses the splatting technique to cast Gaussians within the view frustum onto the camera plane. The projected 2D covariance matrix $\Sigma^{\prime}$ in camera coordinates can be computed as $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$, where $W$ and $J$ denote the viewing transformation and the Jacobian of the affine approximation of the projective transformation, respectively. The color of each pixel $\mathbf{p}$ is calculated by alpha-blending $N$ ordered Gaussians $\{G_{i}|i=1,\cdots,N\}$ overlapping $\mathbf{p}$ as:
$\mathcal{C}=\sum_{i\in N}T_{i}\alpha_{i}\boldsymbol{c}_{i}\text{ with }T_{i}=\displaystyle\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right),$ (6)
where $\alpha_{i}$ denotes the alpha value calculated by multiplying the opacity $o$ with the 2D Gaussian contribution value derived from $\Sigma^{\prime}$, and $T_{i}$ denotes the accumulated transmittance along the ray.
Specifically, each Gaussian’s color is defined by spherical harmonics (SH) coefficients $\mathcal{C}\in\mathbb{R}^{k}$ (where $k$ represents the degrees of freedom). While high-order SH coefficients can contribute to the overall lighting of a scene under various view directions, they are not specifically designed to handle view-dependent color directly, such as specular reflections and shadows.
D.2 Blinn-Phong Reflection Model
The Phong model is widely adopted in computer graphics for simulating the interaction of light with object surfaces [18]. In this model, the reflected light transport on a surface has three components: ambient ($L_{a}$), diffuse ($L_{d}$), and specular reflections ($L_{s}$). The ambient reflection component represents the constant illumination present in the environment, simulating how light scatters and reflects off other surfaces to establish a baseline brightness level. The diffuse reflection component, governed by Lambertian diffusion law, illustrates the scattering of light in multiple directions upon striking a rough surface, leading to a matte appearance. This component is calculated based on the angle between the light direction ($\mathbf{l}$) and the surface normal ($\mathbf{n}$), ensuring that surfaces facing the light source appear brighter. The specular term is computed based on the angle between the light direction vector and the bisector ($\mathbf{h}$) between the view direction ($\mathbf{v}$) and the light direction ($\mathbf{l}$). The Phong model also incorporates a shininess exponent that governs the spread of the specular highlight, allowing the model to represent the surface’s glossiness. We select the Blinn-Phong model [2] to provide the illumination decomposition priors to 3DGS for its simple yet strong physical prior and computation efficiency. With little addition to the Gaussian’s attribute, our model produces realistic lighting effects.
Appendix E Hyperparameters
Table 11 presents the hyperparameters of MetaGS.
Table 11: Hyperparameters of MetaGS.
      Name       Notation       Value
      learning rate       $\beta$       same as 3DGS [9]
      meta-train inner learning rate       $\alpha$       0.01
      meta-train task number       $m$       $5$
      RGB D-SSIM loss scale       $\lambda$       0.2
      sparse loss scale       $\lambda_{\text{sparse}}$       0.001
      normal prediction loss scale       $\lambda_{\text{n}}$       0.2
      normal residual loss scale       $\lambda_{\text{res}}$       0.001
      flatten loss scale       $\lambda_{\text{s}}$       0.001